# Spin Pumping into two-dimensional systems

### v1

@@ Abstract
In this review, we present recent theoretical developments on spin transport phenomena probed by ferromagnetic resonance (FMR) modulation in two-dimensional systems coupled to magnetic materials. We first address FMR linewidth enhancements induced by spin pumping at interfaces, emphasizing their potential as sensitive probes of superconducting pairing symmetries in two-dimensional superconductors. We then examine FMR modulation due to spin pumping into two-dimensional electron gases formed in semiconductor heterostructures, where the interplay of Rashba and Dresselhaus spin–orbit interactions enables gate-controlled spin transport and persistent spin textures. Finally, we investigate spin pumping in monolayer transition-metal dichalcogenides, where spin-valley coupling and Berry curvature effects lead to valley-selective spin excitations and a spin-current Hall effect. These developments demonstrate that the spin pumping technique provides a versatile tool for probing spin transport and spin-dependent phenomena in low-dimensional systems, offering a basis for future spintronics applications.
  • April 2025
@@ 1 Introduction
Low-dimensional systems have become important platforms for exploring spin transport phenomena. Two-dimensional electron gases (2DEGs) in semiconductor heterostructures, such as GaAs/AlGaAs interfaces [11, 12], provide a controllable environment for studying spin–orbit interactions [13, 14, 15, 16] and their impact on spin transport. Atomically thin materials, such as graphene [17, 18] and transition-metal dichalcogenides (TMDCs) [19], provide versatile platforms, with TMDCs exhibiting strong spin–orbit coupling and broken inversion symmetry [20]. These systems enable access to spin-related phenomena beyond those available in bulk materials, and their compatibility with device integration highlights their potential for scalable spintronics technologies. In this review article, we focus on spin transport in 2DEGs and atomically thin materials, emphasizing the roles of spin–orbit interactions [21, 22], spin–valley coupling [20], and Berry curvature effects [23].
@@
Figure 1: A schematic illustration of spin pumping in a paramagnetic metal/ferromagnet bilayer system. Ferromagnetic resonance is induced by microwave irradiation, leading to the injection of a spin current into the metal via interfacial spin exchange interactions. As a back-action of spin pumping, the spin dynamics of the ferromagnet is modulated. This modulation encodes information about the spin excitation of conduction electrons, allowing the determination of their dynamic spin susceptibility in the adjacent metal. Simultaneously, a magnon carrying angular momentum $-\hbar$ is excited in the magnetic system, flipping the conduction electron spin from +/2Planck-constant-over-$+\hbar/2$ to /2Planck-constant-over-$-\hbar/2$ via interfacial interaction. This process serves as a universal mechanism for the generation of spin currents.
@@ 2.1 Ferromagnetic Resonance
Here, $\gamma_{\rm g}$ ($<0$) is a gyromagnetic ratio and $\alpha_{\rm G}$ is the Gilbert damping constant. For a small amplitude of precessional motion around the $z$-axis ($m_{x},m_{y}\ll m_{z}$), the spin direction oscillates around a stable solution as ${\bm{m}}(t)={\bm{m}}_{0}+\delta{\bm{m}}(t)$. The time-dependent modulation is described by the LLG equation as
@@
$\chi_{\rm FI}$ $=\frac{\hbar\gamma_{\rm g}S_{\rm tot}/\hbar}{\omega-\omega_{\bm{0}}+i\alpha_{\rm G}\omega}.$ (3)
@@ 2.2 Spin mixing conductance
Next, let us consider a junction composed of a ferromagnet and a normal metal (NM) (see Fig. 1(b)). In addition to the Gilbert damping, the interfacial exchange coupling produces additional damping due to spin transfer into the NM. This damping can be described by adding an additional damping torque ${\bm{\tau}}$ to the right-hand side of the LLG equation (1): [^1] The torque in the direction of ${\bm{m}}\times(d{\bm{m}}/dt)$ is called the damping torque. In some junction systems, a torque in the direction of $d{\bm{m}}/dt$, which is called the field-like torque, becomes relevant. Since our focus is on a microscopic description of spin pumping, we avoid introducing the field-like torque, which typically arise only in phenomenological treatments.
@@
Here, $\delta\alpha_{\rm G}$ ($>0$) is a modulation of the Gilbert damping. The damping torque ${\bm{\tau}}$, which is equal to the loss of angular momentum per unit time per unit volume, is related to the spin current density across the junction, ${\bm{j}}_{s}$, as ${\bm{\tau}}=-{\bm{j}}_{s}/d$ where $d$ is the thickness of the ferromagnet.
@@
Here, $g^{\uparrow\downarrow}_{r}$ is a parameter called the spin mixing conductance [25, 26], with a dimension of the inverse of the area, which represents the efficiency of spin injection into the NM. This method for spin current generation (or spin injection) driven by the continuous spin precession under microwave is called “spin pumping”. Combining Eqs. (4) and (5), we obtain $\delta\alpha_{\rm G}=g^{\uparrow\downarrow}_{r}/(4\pi\tilde{S}d)$, where $\tilde{S}$ is a spin per unit volume in the ferromagnet. This indicates that a thin ferromagnet is favorable for obtaining a large signal of spin transfer damping.
@@ 2.3 Microscopic description
where $\mathcal{J}>0$ is the exchange coupling constant, $\langle n,m\rangle$ represents summation over all the nearest-neighbor sites, ${\bf\it S}_{n}$ is the localized spin at site $n$ in the FI, and $h_{\mathrm{dc}}$ is the static external magnetic field. We define the Fourier transformation of the spin ladder operators, defined as $S^{+}_{n}=(N_{\rm FI})^{-1/2}\sum_{\bm{k}}S^{+}_{\bm{k}}e^{i{\bm{k}}\cdot{\bm{r}}_{n}}$ and $S^{-}_{\bm{k}}=(S^{+}_{\bm{k}})^{\dagger}$, where $N_{\rm FI}$ is a number of the localized spins in the FI, $S_{n}^{+}=S_{n}^{x}+iS_{n}^{y}$, and ${\bm{r}}_{n}$ is a position of the site $n$. Using the Holstein-Primakoff transformation [38], these ladder operators are expressed with magnon annihilation and creation operators as
@@
[…]
@@ 2.4 Enhanced Gilbert damping
Here we have assumed that a FMR peak is sharp enough ($\alpha_{\rm G}+\delta\alpha_{\rm G}\ll 1$). The present microscopic calculation reveals the physical interpretation of the enhancement of the Gilbert damping: $\delta\alpha_{\rm G}$ is related to the imaginary part of the uniform (local) spin susceptibility of the target material for a clean (dirty) interface.  [^3] We note that the real part of the self-energy has been absorbed into the FMR frequency $\omega_{\bm{0}}$ for simplicity. However, the shift of the FMR frequency also includes information on the target spin susceptiblity and is interpreted in part as the effect of the field-like torque in the context of the spin mixing conductance. For a detailed discussion, see, e.g, Refs. [44, 47]. Our interface model and the corresponding expressions of $J_{{\bf\it k},{\bf\it q}}$ and $\delta\alpha_{\mathrm{G}}$ are summarized in Table 1.
@@
$\chi^{R}({\bm{q}},\omega)=\sum_{{\bm{k}}}\frac{f(\xi_{{\bm{k}}+{\bm{q}}})-f(\xi_{{\bm{k}}})}{\omega+\xi_{\bm{k}}-\xi_{{\bm{k}}+{\bm{q}}}+i\delta}.$ (30)
@@ 2.5 Spin current generation
We calculate the statistical average of the spin current at the interface and treat $H_{\mathrm{T}}$ as a perturbation and $H_{\mathrm{FI}}+H_{\mathrm{sys}}+H_{\mathrm{Z}}$ as an unperturbed Hamiltonian. The second-order perturbation calculation with respect to $H_{\mathrm{T}}$ gives the statistical average of the spin current at the interface [48, 36, 37, 49]
@@
for a clean interface, while $\chi^{R}({\bf\it 0},\omega)$ is replaced with $\sum_{{\bm{q}}}\chi^{R}({\bf\it q},\omega)$ for a dirty interface. For a detail of the calculation, see B. Here, ${\cal A}$ is the interface area and we have assumed that the FMR peak is sharp enough ($\alpha_{\mathrm{G}}+\delta\alpha_{\mathrm{G}}\ll 1$). From this expression, one can see that the interface spin current is characterized by the dynamic spin susceptibility of the adjacent paramagnetic material 444 We note that the nonequilibirum fluctuation of the spin current induced by the spin pumping includes information of a unit of the angular momentum carried by magnetic excitations [50, 51, 52, 53]..
The spin current induced by spin pumping can be detected using the inverse spin Hall effect (ISHE) [24, 54, 55, 56, 27], by which the spin current is converted into a charge current as ${\bm{j}}_{\rm c}=\theta_{\rm SH}(e/\hbar){\bm{j}}_{s}\times\hat{\bm{n}}$ via the spin-orbit interaction in the target material, where $\theta_{\rm SH}$ is the spin Hall angle and $\hat{\bm{n}}$ is a unit normal vector of the interface. To treat such spin-charge conversion, we need to combine another theoretical method for describing electric states in the target material. In Sec. 4 and Sec. 5, we treat the inverse Rashba-Edelstein effect and the valley current, using the Boltzmann equation and the diffusion theory, respectively.
@@ 2.6 Experimental relevance
As mentioned in the Introduction, the spin pumping is a useful experimental tool for obtaining information on spin excitations in two-dimensional systems because of its surface sensitivity. In fact, spin pumping has been observed even for monolayer graphene  [57, 58, 59, 60, 61, 62, 63, 64, 65]. Figure 2 shows an example of the FMR experiment for graphene [58]. The horizontal axis is an external magnetic field, which is proportional to the resonant frequency, while the vertical axis indicates the derivative of the microwave absorption with respect to the magnetic field. We find that the peak of the microwave absorption becomes broad for a permalloy/graphene junction, in comparison with an isolated permalloy. Similar observations have also been reported in a representative TMDC material, $\mathrm{MoS}_{2}$ [66, 67, 68, 69] and two-dimensional Dirac fermions on the surface of three-dimensional topological insulators [70, 71]. These results collectively highlight the potential of FMR as a sensitive spin probe for atomic-scale thin films and surface states, as well as its efficiency as a mechanism for generating spin currents in low-dimensional systems. In the following sections, we explain the microscopic description of spin pumping into specific two-dimensional systems.
Figure 2: Broadening of the FMR linewidth due to spin transfer at the interface. Adapted from Ref [58].
@@ 3.1 Spin-singlet superconductors
[…]
@@ 4 Spin pumping into 2DEG
Semiconductor technology has been an attractive choice for realizing integrated spintronic devices [98, 99]. Two-dimensional electron gas (2DEG) fabricated in semiconductor heterostructures has provided an ideal stage for spintronic devices. The field-effect spin transistor, proposed by Datta and Das [100], is a famous example. In such devices, two kinds of spin-orbit interactions play an important role. One is the Rashba spin-orbit interaction, which appears in systems with broken structural inversion symmetry, mainly induced by confinement potentials in the stacking direction [14, 101]. The other is the Dresselhaus spin-orbit interaction [13], which arises when the crystal structure lacks inversion symmetry, as in zinc-blende III-V and II-IV semiconductors. Both spin-orbit interactions induce spin-momentum locking, i.e., strong dependence of the effective Zeeman field on the electron propagation direction. It is notable that both Rashba and Dresselhaus interactions can coexist and their relative strength can be tuned by a gate voltage [15, 99, 22]. When the strengths of the Rashba and Dresselhaus SOCs are equal, a persistent spin helix (PSH) state emerges [102, 103, 104, 105, 106, 107], leading to prolonged electron spin lifetimes. The Rashba spin-orbit interaction becomes significant also at an interface in the junction systems or heterostructures of transition oxides [108].

### v2

@@ Abstract
In this review, we present recent theoretical developments on spin transport phenomena probed by ferromagnetic resonance (FMR) modulation in two-dimensional systems coupled to magnetic materials. We first address FMR linewidth enhancements induced by spin pumping at interfaces, emphasizing their potential as sensitive probes of superconducting pairing symmetries in two-dimensional superconductors. We then examine FMR modulation due to spin pumping into two-dimensional electron gases formed in semiconductor heterostructures, where the interplay of Rashba and Dresselhaus spin–orbit interactions enables gate-controlled spin transport and persistent spin textures. Finally, we investigate spin pumping in monolayer transition-metal dichalcogenides, where spin-valley coupling and Berry curvature effects lead to valley-selective spin excitations. These developments demonstrate that the spin pumping technique provides a versatile tool for probing spin transport and spin-dependent phenomena in low-dimensional systems, offering a basis for future spintronics applications.
  • October 2025
@@ 1 Introduction
Low-dimensional systems have become important platforms for exploring spin transport phenomena. Two-dimensional superconductors exhibit striking properties such as high transition temperatures and large upper critical fields [11]. The spin transport driven by their distinctive pair potential is therefore of interest. Two-dimensional electron gases (2DEGs) in semiconductor heterostructures, such as GaAs/AlGaAs interfaces [12, 13], provide a controllable environment for studying spin–orbit interactions [14, 15, 16, 17] and their impact on spin transport. Atomically thin materials, such as graphene [18, 19] and transition-metal dichalcogenides (TMDCs) [20], provide versatile platforms, with TMDCs exhibiting strong spin–orbit coupling and broken inversion symmetry [21]. These systems enable access to spin-related phenomena beyond those available in bulk materials, and their compatibility with device integration highlights their potential for scalable spintronics technologies. In this review article, we focus on spin transport in 2DEGs and atomically thin materials, emphasizing the roles of anisotropic superconductivity [22], spin–orbit interactions [23, 24], spin–valley coupling [21], and Berry curvature effects [25].
@@
Figure 1: A schematic illustration of spin pumping in a paramagnetic metal/ferromagnet bilayer system. Ferromagnetic resonance is induced by microwave irradiation, leading to the injection of a spin current into the metal via interfacial spin exchange interactions. As a back-action of spin pumping, the spin dynamics of the ferromagnet is modulated. This modulation encodes information about the spin excitation of conduction electrons, allowing the determination of their dynamic spin susceptibility in the adjacent metal. Simultaneously, a magnon carrying angular momentum -$\hbar$ is excited in the magnetic system, flipping the conduction electron spin from $\hbar/2$ to $\hbar/2$ via interfacial interaction. This process serves as a universal mechanism for the generation of spin currents.
@@ 2.1 Ferromagnetic Resonance (FMR)
Here, $\gamma_{\rm g}$ (<0<0) is a gyromagnetic ratio and $\alpha_{\rm G}$ is the Gilbert damping constant. For a small amplitude of precessional motion around the $z$-axis ($m_{x},m_{y}\ll m_{z}$), the spin direction oscillates around a stable solution as ${\bm{m}}(t)={\bm{m}}_{0}+\delta{\bm{m}}(t)$. The time-dependent modulation is described by the LLG equation as
@@
$\chi_{\rm FI}$ $=\frac{|\gamma_{\rm g}|S_{\rm tot}}{\omega-\omega_{\bm{0}}+i\alpha_{\rm G}\omega}.$ (3)
@@ 2.2 Spin mixing conductance
Next, let us consider a junction composed of a ferromagnet and a normal metal (NM) (see Fig. 1(b)). In addition to the Gilbert damping, the interfacial exchange coupling produces additional damping due to spin transfer into the NM. This damping can be described by adding an additional damping torque $\bm{\tau}$} to the right-hand side of the LLG equation (1): [^1] The torque in the direction of ${\bm{m}}\times(d{\bm{m}}/dt)$ is called the damping torque. In some junction systems, a torque in the direction of $d{\bm{m}}/dt$, which is called the field-like torque, becomes relevant. Although we neglect the field-like torque for simplicity of discussion, we can relate it to the real part of the spin susceptibility, ${\rm Re}\,\chi^{R}({\rm 0},\omega_{\bm{0}})$ [37, 38] (see Sec. 2.4).
@@
Here, $\delta\alpha_{\rm G}$ (>0>0) is a modulation of the Gilbert damping. The damping torque $\bm{\tau}$}, which is equal to the loss of angular momentum per unit time per unit volume, is related to the spin current density across the junction, ${\bm{j}}_{s}$, as ${\bm{\tau}}=-{\bm{j}}_{s}/(\hbar\tilde{S}d)$, where $\tilde{S}$ is a spin per unit volume in the ferromagnet and $d$ is the thickness of the ferromagnet.
@@
Here, $g^{\uparrow\downarrow}_{r}$ is a parameter called the spin mixing conductance [27, 28], with a dimension of the inverse of the area, which represents the efficiency of spin injection into the NM. This method for spin current generation (or spin injection) driven by the continuous spin precession under microwave is called “spin pumping”. Combining Eqs. (4) and (5), we obtain $\delta\alpha_{\rm G}=g^{\uparrow\downarrow}_{r}/(4\pi\tilde{S}d)$. This indicates that a thin ferromagnet is favorable for obtaining a large signal of spin transfer damping.
@@ 2.3 Microscopic description
[…]
@@
$\text{Clean interface:}\quad J_{{\bf\it k},{\bf\it q}}=J\delta_{{\bf\it k},{\bf\it q}},$ (18)
$\text{Dirty interface:}\quad J_{{\bf\it k},{\bf\it q}}=J.$ (19)
@@ 2.4 Enhanced Gilbert damping
Here we have assumed that a FMR peak is sharp enough ($\alpha_{\rm G}+\delta\alpha_{\rm G}\ll 1$). We emphasize that the bulk Gilbert damping, $\alpha_{\rm G}$, is treated phenomenologically, whereas the modulation of the bulk Gilbert damping, $\delta\alpha_{\rm G}$, can be calculated from our microscopic model. The present microscopic calculation reveals the physical interpretation of the enhancement of the Gilbert damping: $\delta\alpha_{\rm G}$ is related to the imaginary part of the uniform (local) spin susceptibility of the target material for a clean (dirty) interface.  [^3] We note that the real part of the self-energy has been absorbed into the FMR frequency $\omega_{\bm{0}}$ for simplicity. However, the shift of the FMR frequency also includes information on the target spin susceptibility and is interpreted in part as the effect of the field-like torque in the context of the spin mixing conductance. For a detailed discussion, see, e.g, Refs. [48, 37]. Our interface model and the corresponding expressions of $J_{{\bf\it k},{\bf\it q}}$ and $\delta\alpha_{\mathrm{G}}$ are summarized in Table 1.
@@
$\chi^{R}({\bm{q}},\omega)=\sum_{{\bm{k}}}\frac{f(\xi_{{\bm{k}}+{\bm{q}}})-f(\xi_{{\bm{k}}})}{\hbar\omega+\xi_{\bm{k}}-\xi_{{\bm{k}}+{\bm{q}}}+i\delta}.$ (30)
@@ 2.5 Spin current generation
We calculate the statistical average of the spin current at the interface and treat $H_{\mathrm{T}}$ as a perturbation and $H_{\mathrm{FI}}+H_{\mathrm{sys}}+H_{\mathrm{Z}}$ as an unperturbed Hamiltonian. The second-order perturbation calculation with respect to $H_{\mathrm{T}}$ gives the statistical average of the spin current at the interface [39, 51, 40, 41, 52]
@@
for a clean interface, while $\chi^{R}({\bf\it 0},\omega)$ is replaced with $\sum_{{\bm{q}}}\chi^{R}({\bf\it q},\omega)$ for a dirty interface. For a detail of the calculation, see B. Here, $\cal A$} is the interface area and we have assumed that the FMR peak is sharp enough ($\alpha_{\mathrm{G}}+\delta\alpha_{\mathrm{G}}\ll 1$). From this expression, one can see that the interface spin current is characterized by the dynamic spin susceptibility of the adjacent paramagnetic material  [^4] We note that the nonequilibrium fluctuation of the spin current induced by the spin pumping includes information of a unit of the angular momentum carried by magnetic excitations [53, 54, 55, 56]..
The spin current induced by spin pumping can be detected using the inverse spin Hall effect (ISHE) [26, 57, 58, 59, 29], by which the spin current is converted into a charge current as $^{\bm{j}}_{\rm c}=\theta_{\rm SH}(e/\hbar){\bm{j}}_{s}\times\hat{\bm{n}}$ via the spin-orbit interaction in the target material, where $\theta_{\rm SH}$ is the spin Hall angle and $\hat{\bm{n}}$ is a unit normal vector of the interface. To treat such spin-charge conversion, we need to combine another theoretical method for describing electric states in the target material. In Sec. 4, we treat the inverse Rashba–Edelstein effect (IREE), using the Boltzmann equation.
@@ 2.6 Experimental relevance
This subsection briefly summarizes the experimental reports on spin pumping into two-dimensional systems. As mentioned in the Introduction, spin pumping is a useful experimental tool for obtaining information on spin excitations in two-dimensional systems, because of its surface sensitivity. Most of the experimental analyses reported so far have relied on the spin-mixing conductance, which does not directly reveal the connection to the underlying microscopic electronic states, while useful for quantifying spin transfer across interfaces. Keeping this in mind, we first introduce several experiments on graphene and outline their development. We then describe recent progress in experiments on TMDC materials, and finally comment on reports involving the surface states of three-dimensional topological insulators.
The FMR damping enhancement induced by spin pumping has been reported for permalloy (Py)/graphene junction systems [60, 61, 62]. Figure 2 (a) shows an example of the FMR experiment for Py/graphene [61]. The horizontal axis is an external magnetic field, while the vertical axis indicates the derivative of the microwave absorption with respect to the magnetic field. We note that the origin of the horizontal axis is chosen as the magnetic field for the FMR resonance, $H_{\rm FMR}$. We find that the peak of the microwave absorption becomes broader for a Py/graphene junction compared with an isolated Py. In addition to the FMR damping enhancement, the same group also observed an inverse spin Hall voltage generated by the injected spin current and reported this as evidence for spin injection into graphene. However, it should be noted that when discussing damping enhancement attributed to spin pumping, structural modifications of the ferromagnet induced by forming the heterostructure can also give the appearance of increased damping [63], and the experimental results should therefore be interpreted with particular caution. Since these pioneering spin pumping experiments on graphene, various experimental groups have reported spin pumping related phenomena, and it is becoming a consensus that spin pumping into graphene is possible and can be characterized by FMR modulation [64, 65, 66, 67, 68, 69, 70].
Similar experimental observations have also been reported by several experimental groups in TMDC materials [71, 72, 73, 74, 75, 76, 77]. Figure 2 (b) shows the effective damping constant of a WS2/Co3FeB (FM) heterojunction plotted as a function of the inverse FM thickness $1/t_{\mathrm{FM}}$. From this result, the authors argued that the FM exhibits a damping enhancement when interfaced with WS2, and that the proportionality to $1/t_{\mathrm{FM}}$ is consistent with spin pumping in the heterojunction. On the other hand, a very recent study reports that in an atomically flat MoSe2/CoFeB heterojunction, no damping enhancement attributable to spin pumping was detected [78]. Careful analysis of the experiments, together with a microscopic understanding of the interfacial electronic states, will be essential for achieving a consistent interpretation of these findings.
Spin transport via two-dimensional Dirac fermions on the surfaces of three dimensional topological insulators has been explored using ferromagnet and topological insulator heterostructures [79, 80, 81, 82, 83]. In these systems, the spin momentum locked surface states enable efficient spin to charge conversion and modify the magnetization dynamics of the adjacent magnetic layer. Spin pumping into these surface states has been reported to cause a noticeable enhancement of the effective damping, and signatures of inverse Edelstein effect and spin torque generation have also been observed. These results indicate that topological insulator surfaces offer a promising platform for spin current generation and detection.
Experimental studies on graphene, TMDCs, and three-dimensional topological insulators have advanced considerably, and the description in terms of spin mixing conductance has played an important role in quantifying spin transfer in the early stages. This approach, however, does not directly reveal the connection to characteristic microscopic electronic structures such as massless Dirac fermions, spin valley coupling, or spin momentum locking. To gain deeper insight, it is important to develop and apply microscopic theoretical descriptions alongside experimental progress so that both can advance together. The following sections provide such a theoretical framework for spin pumping into specific two-dimensional systems.
Figure 2: (a) Broadening of the FMR linewidth due to spin transfer at the interface between 25-nm-thick permalloy (Py) and single-layer graphene. Measurements were performed at room temperature and microwave frequency $9.62~\mathrm{GHz}$. Adapted from Ref. [61]. (b) Dependence of the effective damping enhancement on the Co3FeB thickness ($t_{\mathrm{FM}}$) in a WS2/Co3FeB heterostructure. Adapted from Ref. [75].
@@ 3.1 Spin-singlet superconductors
$\delta\alpha_{\mathrm{G}}=\frac{2\pi S_{0}J^{2}D_{\mathrm{n}}(\epsilon_{\rm F})^{2}}{\hbar\omega}\int_{-\infty}^{\infty}dE\,[f(E)-f(E+\hbar\omega)]F(E,\omega)D(E)D(E+\hbar\omega),$ (41)
@@ 4 Spin pumping into 2DEG
Semiconductor technology has been an attractive choice for realizing integrated spintronic devices [108, 109]. Two-dimensional electron gas (2DEG) fabricated in semiconductor heterostructures has provided an ideal stage for spintronic devices. The field-effect spin transistor, proposed by Datta and Das [110], is a famous example. In such devices, two kinds of spin-orbit interactions play an important role. One is the Rashba spin-orbit interaction, which appears in systems with broken structural inversion symmetry, mainly induced by confinement potentials in the stacking direction [15, 111]. The other is the Dresselhaus spin-orbit interaction [14], which arises when the crystal structure lacks inversion symmetry, as in zinc-blende III-V and II-VI semiconductors. Both spin-orbit interactions induce spin-momentum locking, i.e., strong dependence of the effective Zeeman field on the electron propagation direction. It is notable that both Rashba and Dresselhaus interactions can coexist and their relative strength can be tuned by a gate voltage [16, 109, 24]. When the strengths of the Rashba and Dresselhaus SOCs are equal, a persistent spin helix (PSH) state emerges [112, 113, 114, 115, 116, 117], leading to prolonged electron spin lifetimes. The Rashba spin-orbit interaction becomes significant also at an interface in the junction systems or heterostructures of transition oxides [118].